Let V be a Hilbert space. Let S1 and S2 be two hyperplanes in V defined by

S_1 = \left \{ x\in V |\left \langle a_1,x \right \rangle = b_1 \right \},S_2 = \left \{ x\in V |\left \langle a_2,x \right \rangle = b_2 \right \}

Let y\in V be given. We consider the projection of y onto Sin S2, i.e., the solution of

min TESInS2 (1)

(a) Prove that Sin S2 is a plane, i.e., if x,z \in S_1 \cap S_2 , then (1+t)z - tx \in S_1 \cap S_2 for any t\in\mathbb{R}.

(b) Prove that z is a solution of (1) if and only if z \in S_1 \cap S_2 and

\left \langle z-y,z-x \right \rangle = 0,\forall x \in S_1 \cap S_2 (2)

(c) Find an explicit solution of (1). (

d) Prove the solution found in part (c) is unique.



Sin S2
min TESInS2
Answer #1

a) We know x,z\in S_1\cap S_2 if and only if

\langle a_1,x\rangle=b_1=\langle a_1,z\rangle,~~~~\langle a_2,x\rangle=b_2=\langle a_2,z\rangle

Thus, for any t\in\mathbb R we have

a1 .tx + (1 _ t)z)-t(a1·z) + (1 _ t)(a1·z) thị + (1 _ t)b1-b1

and

\langle a_2,tx+(1-t)z\rangle=t\langle a_2,x\rangle+(1-t)\langle a_2,z\rangle=tb_2+(1-t)b_2=b_2

This shows that tx+(1-t)z\in S_1\cap S_2 . Hence, Sin S2 is a plane.

b) Let z be a solution. Then

||z-y||=\min_{x\in S_1\cap S_2}||x-y||

shows that it is in Sin S2 because \begin{align*}S_1,S_2\end{align*} are closed. (Alternatively, otherwise, the "line segment" \begin{align*}tz+(1-t)y\end{align*} will intersect at a point which will make \begin{align*}||x-y||\end{align*} smaller). Also, for any x\in S_1\cap S_2 let \begin{align*}\alpha=\langle z-y,z-x\rangle\end{align*} . Suppose that \begin{align*}\alpha\neq 0\end{align*} for some \begin{align*}x\end{align*} in Sin S2, and without loss of generality let us assume \begin{align*}||z-x||=1\end{align*} ; then we have

\begin{align*}||y-(\alpha x+(1-\alpha)z||^2&=||(y-z)-\alpha(x-z)||^2\\ &=||y-z||^2+\alpha^2||x-z||^2-2\alpha\langle z-y,z-x\rangle\\ &=||y-z||^2+\alpha^2||x-z||^2-2\alpha^2\\ &=||y-z||^2-\alpha^2\\ &<||y-z||^2\end{align*}

a contradiction to

||z-y||=\min_{x\in S_1\cap S_2}||x-y||

since \begin{align*}\alpha=\langle z-y,z-x\rangle\end{align*} is zero.

Because this argument is reversible, we get the only if part via backtracking.